# Kaiwu: A Multimodal Manipulation Dataset and Framework for Robot Learning and Human-Robot Interaction

### v1

@@ II-A Robot learning dataset
Multi-modal input can serve as a crucial foundation for robots to learn human behavior, enhancing the accuracy and efficiency of the learning process. The dataset focuses on teaching robots to perform daily interactions in changing environments through human demonstrations [7]. JD ManiData, ManiWAV [11] and RH20T [12] are designed to facilitate the acquisition of diverse and generalizable skills by robots, and the robot hardware platform for DROID data [13] is collected through a distributed collection approach to improve the performance and generalization of the robot’s operational strategies. A large-scale benchmark dataset called GraspNet-1Billion is proposed by Fang et al.[14], aiming to improve robust object grasping in cluttered scenes under multimodal data conditions. The Open-X implementation [15] enables the integration of large-scale data volumes and validates the positive migration of data between robots to improve multi-robot capabilities. However the above datasets suffer from data homogenization. The ability to efficiently and purposefully train embodied intelligence models in complex scenarios is limited. Building on the excellent datasets above, more research has been performed to further extend on a particular aspect. The evolution of the RT series datasets exemplifies this progression. The Robotics Transformer 1 (RT-1) [8] is designed to achieve multi-task robot control in real-world environments by utilizing large-scale datasets and high-capacity model architectures. The RT-1 model effectively generalizes by incorporating datasets that encompass a wide range of robotic operations and environmental scenarios. The RT-1 model dataset comprises text instructions, robotic vision images, image frame actions, and commands. However, there are limitations in the selection of scenes and target objects. In the subsequent RT Series models, the datasets are continuously expanded in actions, target objects and target position aspects. RT-2 [9], building on the foundation of RT-1, utilizes Web VQA (Web-based Visual Question Answering) to match the image frames from RT-1 with corresponding text tokens, effectively addressing the limitations related to target objects and positions. The Open X-Embodiment dataset [15] proposes the integration of the RT-2 database with additional datasets containing various actions and skills. The initiative aims to train the RT-X model, which is designed to excel in versatility across diverse usage scenarios. In the latest RT-H, the dataset has been augmented with human-robot interaction data pertaining to human interventions, thereby enhancing the model’s learning capabilities. In addition, the ARIO [16] dataset is optimized based on the Open X-Embodiment dataset. It contains more sensory data in 5 modalities: 2D, 3D, text, tactile, and sound, and covers two types of tasks: operation and navigation. It contains both simulation data and real scene data, and includes a variety of robot hardware subjects, retaining a large amount of data and a uniform format. However, These datasets are limited by insufficient dynamic kinematic information and a lack of authenticity in the raw data, and suffer from the issue of labeling generality.
@@ II-B Human activity recognition dataset
Enabling robots to accurately recognize and understand human activity is an essential aspect of developing embodied intelligence. Therefore, high-quality datasets are urgently needed. The industry-oriented dataset [17] is a dataset of human movement in industrial activities and is intended to be used to classify, predict or evaluate human movement in an industrial environment. The ActionSense dataset proposed by DelPreto et al. [18], which focuses on kitchen scenes, provides a multimodal human activity dataset. Yoon et al. [19] proposed HUMBI which is a multiview camera dataset for human body modeling. The dataset addresses the limitations of human expression datasets by modeling human appearance in high resolution. The HUMBI dataset is not only significant for appearance modeling but also serves as a powerful multi-view benchmark for establishing more robust and generalizable human body reconstruction and recognition models. This, in turn, enhances the identification capabilities of agents. The Toyota Smarthome Untrimmed (TSU) dataset proposed by Dai et al. [20] collects untrimmed video data of spontaneous activities in home settings. This untrimmed data can help robots understand the causality of human activities before and after, thereby enhancing their ability to recognize human behaviors. The dataset [21] is designed to help autonomous robots become aware of the overall intent of the team they are working with in order to effectively help their human counterparts or augment the team’s capabilities. The current research lacks an extension of the application scenarios for the dataset and a focus on assembly task related scenarios. In addition, there is a lack of balance on data modality homogeneity, cross-modality temporal consistency, and task causality analysis.
@@ II-C Human robot collaboration dataset
Embodied intelligence emphasizes the interaction capabilities of intelligent systems with their physical bodies and environments. Human-robot collaboration enables agents to learn how to better adapt to and comprehend their surroundings through interactions with humans, thereby enhancing the generalizability and stability of behavioral patterns. Multimodal datasets integrate various data types, such as text, images, sound, and video, which contain rich information about user intentions and emotions. In human-robot interaction, multimodal data can assist agents in better understanding the needs of operators, improving accessibility and information perception capabilities, and providing a more natural and intelligent mode of interaction. The dataset [22] focuses on teaching robots to perform daily interactions in changing environments through human demonstrations. The RBO dataset [23] records data on human interaction with articulated objects, which provides a valuable data resource for robotics and inspires the construction of subsequent human-robot interaction datasets. The RT-H [10] method adds interventions of human-robot interaction behaviors to the RT-2, which enhance the robot’s understanding of interaction action or language by describing motions with more fine-grained phrases. Thereby, RT-H builds an action hierarchy, enabling the agent to learn from the interventions of human language motions.
Additionally, eye gaze, EMG signals, IMU and skeletal postures contain a wealth of information about human operational intent that can be utilized to enhance the accuracy and diversity of collaborative human-robot interactions. Mapping the region of interest using eye-tracking devices along the viewing trajectory can help the robot understand the human’s intentions during the assembly process. EMG and IMU signals record the behavior patterns of human arm muscles during the execution of an action. This data can help the robot learn and predict hand gestures to respond to human behavior. The HARMONIC dataset [24] constitutes a comprehensive collection of human-robot interactions within a shared autonomous setting. Throughout the data acquisition process, the robot executes a series of experimental maneuvers facilitated by a joystick which is controlled by a person with wearable sensors. This approach enables the aggregation of data that is rich in human intent, thereby providing valuable insights into the nuances of human-robot interaction. Consequently, the HBOD dataset [kang2023hbod] employs more wearable motion sensors to capture intricate details of human movement and direct interactions with an array of tools. This methodology significantly enhances the robot’s capacity to comprehend human movement intentions and strategic maneuvers within a specified context. The dataset [25] facilitates an expansive discourse on overarching learning challenges within the realm of hand-object interaction, encompassing aspects such as intention recognition and motion generation, with implications for the field of Robotics. The OAKINK2 dataset [26] provides multi-view image streams and accurate pose annotations of the human body, hands, and various interacting objects. This rich data supports a variety of applications such as interaction reconstruction and motion synthesis.
The aforementioned datasets employ indirect methods of data collection, which leads to deficiencies in recording and describing dynamic information. The action interactions within the designed scenarios fail to connect motion units into a narrative process, lacking in causality and coherence. In Kaiwu dataset, wearable sensors are utilized to collect dynamic and static data directly. The scenario design integrates assembly action units into the entire assembly process, rendering the human-robot interaction more narrative and causally connected. Furthermore, the interaction process is complemented with the establishment of a coordinate framework for experimental environment, thereby enriching the dynamic information and spatial integrity of the interaction. A detailed comparison of the proposed methods with other SOTA datasets was performed (see Table I).
@@ III Data collection platform setups
The data collection platform facilitates the synchronized streaming, storage, and visualization of info from wearable and environment-mounted sensors with varying sampling rates and data formats. The platform’s structure simplifies the integration of newly added sensors. The multi-threading and multi-processing is designed to ensure efficient CPU and RAM usage during the collection. Additionally, it offers post-processing capabilities that allow for the combination of streamed data with data processing by third-party applications.
The platform also aims to facilitate efficiency of data collection with fewer experimental artifacts by allowing researchers to operate, record metadata without tedious post-processing, and pay more attention to the process of subject build.
TABLE II: Overview of sensors setups.
Device Sensor Type Data Sreams Sampling rate [Hz] Calibration Third-Party Recording Software
Tactile sensors grip force feedback
WISEGLOVE19FE Angle sensors finger angel 100 hand pose calibration GraspMF
Arm IMU hands,arm quaternion
Trigno Biofeedback System EMG , IMU EMG signal,ACC 4,000 Stand with known locations and poses —————————————
First-person camera First-person videos Tobii Lab
Tobii Pro Glasses 3 Infrared eye camera gaze point 50/100 Pupil Capture calibration Tobii Pro Eye Tracker Manager
IMU pupil details
Nokov XINGYING Motion capture camera 340 Stand with known poses and locations XINGYING system 2.1.0
Color videos
Azure Kinect DK RGB+D raw-format frame images 60 Place with participant in field of view —————————————
Depth data
Microphone Omnidirectional,cardioid Raw audio recordings 48,000 Place in preset position —————————————
@@ III-B1 Data glove
TABLE III: Data format of glove data.
@@
Data gloves are utilized to simultaneous capture the hand movement and palm tactile interaction information. Therefore, the WISEGLOVE glove (Fig. 2) is chosen which can drive 3D animated human hand movements in real time. 19 finger angle sensors and 19 finger pressure sensors are equipped with a sensing accuracy of 9g, which fulfills the spatial and accuracy resolution of the experimental requirements. Additionally, there is an arm sensor that recorded quaternion data from the palm, forearm, and upper arm sensors. The system can capture human hand motion in real-time and drives 3D animation simultaneously to depict hand motions. This system is used to study participants’ hand movements and finger coordination during assembling tasks.
@@ III-B2 EMG and ACC
TABLE IV: Data format of EMG (or ACC) data.
@@ III-B3 Environment depth and visual information
Kinect is positioned directly in front of the participant to record the environmental information including RGB and depth information (Fig. 3) during the assembly process. This camera is used to capture participants’ body postures and movements in real-time for investigating their motor performance and movement control in specific tasks.
Figure 3: RGB-D information.
@@ III-B4 Eye tracker
Eye-tracking data, including participants’ gaze points mobility, are obtained to invesitgate visual attention and cognitive processes. The Tobii Eye Tracker (Fig. 4) is a binocular stereoscopic dark pupil tracking device that relies on corneal reflexes to record and analyze human eye tracking and first-person information in assembly processes.With an average accuracy of 0.6°, an average precision of 0.03°and an average data loss rate of 0.01%, the Tobii Pro Glasses 3 guarantees quality data collection.
Figure 4: Attention information of eye tracker.
@@ III-B6 Ground truth
The Nokov Optical 3D Motion Capture System is used to capture and record the movement and actions of a person or object in three-dimensional space. It utilizes high-precision optical sensors and cameras to track the movement trajectory of the object being measured and then transmitting the data to a computer for analysis and rendering. In the experiment, participants are fitted with reflective markers at 37 key points on the body to calibrate and build a mannequin. With this system, the ground truth information for analysis can be accurately captured. The location of reflective feature points, camera placement, and human skeleton modeling of the motion capture system are shown in Fig. 5.
Figure 5: Motion capture system calibration.
@@ III-C1 Equipment initialization calibration
The human eye exhibits natural individual differences in shape and geometry. The location of the visual central fovea, responsible for perceiving the details and colors of a focused target, varies from person to person. To address individual differences in eye tracking performance, the Tobii Pro eye tracker utilizes a one-point calibration procedure to enhance the eye tracker’s accuracy in predicting the subject’s visual position. To ensure that accurate eye-tracking data is obtained in an experiment, the eye-tracker must be calibrated for each subject. The data recorder positions the calibration card facing the subject, ensuring that the card is held at an arm’s length away from the participant. The participant should wear the eye-tracker’s head-mounted module and maintain focus on the center of the calibration card throughout the calibration process.
Physical stature varies across participants, so for each collection experiment, the motion capture system needs to be calibrated to the skeleton of the participants after the reflective markers are installed. The participant’s arms are opened at shoulder level and slowly rotated to ensure that all reflective markers on the participant’s body are recorded in the system (Fig. 5).
@@ IV-A Participants
Twenty participants with a mean height of 172.4$\pm$6.5 cm and a mean age of 23.95$\pm$5.1 yrs are recruited in this process, and all the consents are given. Each volunteer is recorded for 2 hours. The entire data collection process span 11 working days. Approval of all ethical and experimental procedures and protocols is granted by Institutional Review Board of Tongji University under Application No.TJDXSR055, 2023. The experiments are performed in the Kaiwu Lab of Tongji University.
@@ IV-B1 Assembly activity
In the assembly process of the robotic arm, links that closely reflect human dexterity and dynamics are selected for data collection. The entire assembly process is divided into 15 assembly links, with each link divided into left and right fine-grain motion elements labeled as action units, which are recorded in glove exported model playback. The segmentation facilitates the coding and integration of the entire assembly process into an action ensemble, and also facilitates subsequent labeling and recall. The assembly activities are shown in Fig. 6.
@@ IV-B2 Assembly equipment and parts
For the above action set, the relevant set of frontal assembly objects includes the robot front arm, robot end arm, robot middle arm, robot arm base, rear arm part, flange bearings, linkage, motor, steering bearings, brass post, steel post, bolt spacer, bolts, nuts, and power cable.
Instead of subdividing the action units into finer degrees in our experiments, we remain at the level of a relatively large patchwork of components. This is due to the fact that it is difficult for the finger tactile sensors to detect changes of data in small and delicate components. Before proceeding with data collection, an initial pre-assembly of the original small components is conducted to meet the acquisition requirements.
@@ V Data annotation
The annotation information, such as labels and instance, is shown in Table V.
TABLE V: Overview of annotation.
@@ V-A Action&gesture segmentation and export
For action segmentation, the left and right-handed action segments of each participant are divided as a whole into time intervals for the stage actions. And the free gaps between movements are retained to preserve the causality and coherence between movements. The segmentation of movements includes basic and specific movements, with the former including generalized movements such as approaching, grasping, placing, and tightening screws. The latter includes actions specific to each assembly-specific task, such as the wiring action in the motor mounting session.
@@
Firstly, according to whether or not the palm is involved, then they are divided into Group1: Cyl, Obl, IntPP and Group2: Pinch, LatP, SpP, Lum. in Group1 is further divided by the state of the thumb and the four fingers, which corresponds to Cyl: the thumb and the four fingers are naturally curved; Obl: the thumb is straight and the four fingers are curved; and the others are stipulated as IntPP. in Group2 is further divided by the number of fingers involved in the The number of fingers involved in the activity is further classified as Group2a: a small number (2-3) of fingers involved, containing Pinch and LatP; Group2b: four fingers involved, containing Lum and SpP.Finally, by the presence or absence of the lateral fingers, it is classified as LatP and SpP with lateral finger involvement, and Pinch and Lum without lateral finger involvement. The final label distribution is shown in Fig. 8.
@@ V-B Semantic segmentation
Semantic information annotation of 30 key objects in RGB video data is performed. By annotating the environment information from Kinect, which records the third-person view of the participants, the dataset is enriched with semantic information following the assembly logic. This enhancement aims to improve the robot’s logical understanding of the assembly process and human action learning. Labels and instances of annotation are shown in the Fig. 9.
@@ VI-A Descriptive statistics
A total of $6\times 15\times 20$ sets of experimental data on the assembling process are collected for Kaiwu dataset. The data comes from wearable sensors worn by 20 participants. Each participant is asked to complete 15 typical actions specified in the assembly process. Each participant spend approximately 19 minutes, resulting in the dataset representing about 6.3 hours of the assembling process. A summary of the related data is presented in Table VI.
TABLE VI: Descriptive Statistics.
@@ VI-B1 EMG data
The EMGData folder contains the ACC data output by the built-in IMU sensors and the EMG files measured and output by the EMG sensors. The ACC or EMG data are output by each of the built-in IMUs or EMG sensors respectively in 17 sets of data recorded as TIME.csv files. Each file contains 1 row for the calibration standard value and 2-17 rows for the ACC or EMG signals from the 16 sensors. The EMG_data directory structure is shown in Fig. 11.
{forest}
forest scheme [EMG or ACC data [ACC [C1-source_data [TIME.csv] ] [⋮] [C15-source_data] ] [EMG [C1-source_data [TIME.csv] ] [⋮] [C15-source_data] ] ]
Figure 11: Structure of the EMG or ACC data.

### v2

@@ II-A Robot learning dataset
Multi-modal input can serve as a crucial foundation for robots to learn human behavior, enhancing the accuracy and efficiency of the learning process. The dataset focuses on teaching robots to perform daily interactions in changing environments through human demonstrations [7].
JD ManiData, ManiWAV[11], and RH20T [12] help robots acquire diverse and generalizable skills. The DROID dataset [13] uses a distributed approach to enhance robot performance and generalization. GraspNet-1Billion [14] improves object grasping in cluttered scenes with multi-modal data. The Open-X implementation [15] enables large-scale data integration and validates data transfer between robots to improve multi-robot capabilities.
However, the above datasets suffer from data homogenization. The ability to efficiently and purposefully train embodied intelligence models in complex scenarios is limited. Building on the excellent datasets above, more research has been done to further extend on a particular aspect.
The RT series datasets evolve to enhance robot capabilities. RT-1 [8] uses large-scale data and high-capacity models for multitask control in real-world environments, incorporating text, vision, and action data. However, it has limitations in scene and target object selection. RT-2 [9] builds on RT-1 using Web VQA to link images with text tokens, addressing these limitations. The Open X-Embodiment dataset [15] integrates RT-2 with more diverse data to train the versatile RT-X model. The latest RT-H adds human-robot interaction data to further improve learning capabilities.
The ARIO dataset [16], based on Open X-Embodiment, includes simulation and real-world data with diverse robot hardware and a uniform format. However, it lacks dynamic kinematic information and authenticity in raw data, and faces labeling generality issues.
@@ II-B Human activity recognition dataset
Enabling robots to recognize and understand human activity is crucial for embodied intelligence. High-quality datasets are essential. The industry-oriented dataset [17] captures human movement in industrial settings for classification and prediction. ActionSense [18] focuses on kitchen scenes with multi-modal data. HUMBI [19] provides a multiview camera dataset for high-resolution human body modeling, enhancing reconstruction and recognition capabilities. The Toyota Smarthome Untrimmed (TSU) dataset [20] collects untrimmed home videos to help robots understand the causality of human activity. Another data set [21] aims to help robots understand the intention of the team to better assist humans.
The current research lacks an extension of the application scenarios for the dataset and a focus on assembly task-related scenarios. In addition, there is a lack of balance in data modality homogeneity, cross-modality temporal consistency, and task causality analysis.
@@ II-C Human robot collaboration dataset
Embodied intelligence focuses on how intelligent systems interact with their physical bodies and environments. Human-robot collaboration helps agents adapt to surroundings by learning from human interactions, improving the generalizability and stability of their behavior. Multimodal datasets, which include text, images, sound, and video, provide rich information about user intentions and emotions. In human-robot interaction, these datasets enhance agents’ understanding of user needs, improve accessibility and perception, and enable more natural and intelligent interactions.
The dataset [22] teaches robots daily interactions in changing environments via human demonstrations. The RBO dataset [23] records human interactions with articulated objects. The RT-H [10] method enhances RT-2 by adding human-robot interaction interventions, using fine-grained phrases to describe motions. This improves the robot’s understanding of interaction actions and language, builds an action hierarchy, and enables learning from human language interventions.
Eye-tracking devices can map the region of interest along the viewing trajectory, helping robots understand human intentions during tasks like assembly. EMG and IMU signals capture muscle behavior patterns during actions, enabling robots to learn and predict hand gestures and respond more accurately to human behavior.
The HARMONIC dataset [24] captures human-robot interactions in an autonomous setting, using a joystick-controlled robot and wearable sensors to gather data rich in human intent. The HBOD dataset [25] employs more wearable sensors to capture detailed human movements and interactions with tools, enhancing robot understanding of human intentions and maneuvers. The dataset [26]explores learning challenges in hand-object interactions, including intention recognition and motion generation. The OAKINK2 dataset [27] provides multi-view images and accurate pose annotations of humans and objects, supporting applications like interaction reconstruction and motion synthesis.
Regarding some shortcomings of the above datasets,the Kaiwu dataset directly collects dynamic and static data using wearable sensors. It integrates assembly actions into a coherent process, enhancing human-robot interaction with a narrative and causal structure. Additionally, a coordinate framework is established to enrich spatial integrity. A detailed comparison with other state-of-the-art datasets is provided in Table I.
@@ III Data collection platform setups
The data collection platform enables synchronized streaming, storage, and visualization of information from wearable and environment-mounted sensors, accommodating different sampling rates and data formats. Its modular structure simplifies the integration of new sensors. Multi-threading and multi-processing ensure efficient CPU and RAM usage. The platform also supports post-processing, allowing integration with third-party applications. It aims to enhance data collection efficiency by reducing experimental artifacts, enabling easy metadata recording, and allowing researchers to focus more on the subject build process.
@@ III-B1 Data glove
TABLE II: Data format of glove data.
@@
Data gloves are utilized to simultaneous capture the hand movement and palm tactile interaction information. Therefore, the glove (Fig. 2) is chosen which can drive 3D animated human hand movements in real time. 19 finger angle sensors and 19 finger pressure sensors are equipped with a sensing accuracy of 9g, which fulfills the spatial and accuracy resolution of the experimental requirements. Additionally, there is an arm sensor that recorded quaternion data from the palm, forearm, and upper arm sensors. The system can capture human hand motion in real-time and drives 3D animation simultaneously to depict hand motions. This system is used to study participants’ hand movements and finger coordination during assembling tasks.
@@ III-B2 EMG and ACC
TABLE III: Data format of EMG (or ACC) data.
@@ III-B3 Environment depth and visual information
Camera is positioned directly in front of the participant to record the environmental information including RGB and depth information during the assembly process. This camera is used to capture participants’ body postures and movements in real-time for investigating their motor performance and movement control in specific tasks.
@@ III-B4 Eye tracker
Eye-tracking data, including participants’ gaze points mobility, are obtained to invesitgate visual attention and cognitive processes. The eye tracker is a binocular stereoscopic dark pupil tracking device that relies on corneal reflexes to record and analyze human eye tracking and first-person information in assembly processes. With an average accuracy of 0.6°, an average precision of 0.03°and an average data loss rate of 0.01%, the eye tracker device guarantees quality data collection.
@@ III-B6 Ground truth
The 3D Motion Capture System is used to capture and record the movement and actions of a person or object in three-dimensional space. It utilizes high-precision optical sensors and cameras to track the movement trajectory of the object being measured and then transmitting the data to a computer for analysis and rendering. In the experiment, participants are fitted with reflective markers at 37 key points on the body to calibrate and build a mannequin. With this system, the ground truth information for analysis can be accurately captured.
@@ III-C1 Equipment initialization calibration
The eye tracker uses a one-point calibration procedure to account for individual differences in eye shape and geometry, enhancing accuracy in predicting visual positions. Before experiments, the eye tracker is calibrated for each subject. The data recorder positions a calibration card at arm’s length from the subject, who wears the eye tracker’s head-mounted module and focuses on the card’s center during calibration.
Physical stature varies across participants, so for each collection experiment, the motion capture system needs to be calibrated to the skeleton of the participants after the reflective markers are installed. The participant’s arms are opened at shoulder level and slowly rotated to ensure that all reflective markers on the participant’s body are recorded in the system.
@@ IV-A Participants
Twenty participants with a mean height of 172.4$\pm$6.5 cm and a mean age of 23.95$\pm$5.1 yrs are recruited in this process, and all the consents are given. Each volunteer is recorded for 2 hours. The entire data collection process span 11 working days. Approval of all ethical and experimental procedures and protocols is granted by Institutional Review Board.
@@ IV-B1 Assembly activity
In the assembly process of the robotic arm, links that closely reflect human dexterity and dynamics are selected for data collection. The entire assembly process is divided into 15 assembly links, with each link divided into left and right fine-grain motion elements labeled as action units. The segmentation facilitates the coding and integration of the entire assembly process into an action ensemble, and also facilitates subsequent labeling and recall. The assembly activities are shown in Fig. 3.
@@ IV-B2 Assembly equipment and parts
For the action set, the relevant assembly objects include the robot front arm, end arm, middle arm, arm base, rear arm part, flange bearings, linkage, motor, steering bearings, brass post, steel post, bolt spacer, bolts, nuts, and power cable.
In our experiments, we do not subdivide the action units into finer degrees. Instead, we work with larger components because finger tactile sensors struggle to detect data changes in small, delicate parts. Before data collection, an initial pre-assembly of the small components is performed to meet acquisition requirements.
@@ V Data annotation
The annotation information, such as labels and instance, is shown in Table IV.
TABLE IV: Overview of annotation.
@@ V-A Action&gesture segmentation and export
For action segmentation, the left and right-handed action segments of each participant are divided as a whole into time intervals for the stage actions. And the free gaps between movements are retained to preserve the causality and coherence between movements.
@@
Firstly, according to whether or not the palm is involved, then they are divided into Group1: Cyl, Obl, IntPP and Group2: Pinch, LatP, SpP, Lum. in Group1 is further divided by the state of the thumb and the four fingers, which corresponds to Cyl: the thumb and the four fingers are naturally curved; Obl: the thumb is straight and the four fingers are curved; and the others are stipulated as IntPP. in Group2 is further divided by the number of fingers involved in the The number of fingers involved in the activity is further classified as Group2a: a small number (2-3) of fingers involved, containing Pinch and LatP; Group2b: four fingers involved, containing Lum and SpP.Finally, by the presence or absence of the lateral fingers, it is classified as LatP and SpP with lateral finger involvement, and Pinch and Lum without lateral finger involvement.
@@ V-B Semantic segmentation
Semantic information annotation of 30 key objects in RGB video data is performed. By annotating the camera environment information, which records the third-person view of the participants, the data set is enriched with semantic information following the assembly logic. This enhancement aims to improve the robot’s logical understanding of the assembly process and human action learning.
@@ VI-A Descriptive statistics
A total of $6\times 15\times 20$ sets of experimental data on the assembling process are collected for Kaiwu dataset. The data comes from wearable sensors worn by 20 participants. Each participant is asked to complete 15 typical actions specified in the assembly process. Each participant spend approximately 19 minutes, resulting in the dataset representing about 6.3 hours of the assembling process. A summary of the related data is presented in Table V.
TABLE V: Descriptive Statistics.
@@ VI-B1 EMG data
The EMGData folder contains the ACC data output by the built-in IMU sensors and the EMG files measured and output by the EMG sensors. The ACC or EMG data are output by each of the built-in IMUs or EMG sensors respectively in 17 sets of data recorded as TIME.csv files. Each file contains 1 row for the calibration standard value and 2-17 rows for the ACC or EMG signals from the 16 sensors.